OpenStudy (anonymous):

use the formula d=rt to find the value of the missing variable d=130 feet, t=26 seconds r=____ft/sec

OpenStudy (anonymous):

When you plug in these values (130 and 26) into your formula (d=rt), what does the formula look like?

OpenStudy (anonymous):

@BangkokGarrett IS IT 130/26 =5

OpenStudy (texaschic101):

d = rt ....your looking for r so you can change the formula to d/t = r. Now lets do some subbing... d/t = r 130/26 = r 5 = r you are correct.....r = 5

OpenStudy (anonymous):

5 ft/s...yes, you got it!